OpenStudy (anonymous):

slope of (2,16),(-8,-5)

OpenStudy (anonymous):

To find the slope, you should use the formula of slope: y1-y2/x1-x2=slope.

OpenStudy (anonymous):

The slope is 16-(-5)/2-(-8)=21/11.